If a pebble is dropped toward a pond from the top of a 360-ft building, its height, h, after t seconds is given by h = -16t^2 + 180. How long does it take the pebble to pass through a height of 60 feet?
Found 2 solutions by stanbon, hulamc723:
Answer by stanbon(48568)

h = -16t^2 + 180. How long does it take the pebble to pass through a height of 60 feet.
---------------------
60=-16t^2+180
16t^2=120
T^2=15/2
t=(sqrt30)/2 seconds or 2.739 seconds
Cheers,
Stan H.

Answer by hulamc723(15)

If a pebble is dropped toward a pond from the top of a 360-ft building, its height, h, after t seconds is given by h = -16t^2 + 180. How long does it take the pebble to pass through a height of 60 feet?
well, since the height is 60, we can substitute 60 for h. we now have
60=-16t^2+180. Subtract 60 from both sides and you get 0=-16t^2+120 Now use the quadratic formula to solve (x= -b +or- the square root of b^2-4(a)(c)/2a)
In this equation a=-16, b=0, and c=120. So, now we have x=0 + or - the square root of 0-4(-16)(120)/2(-16) or x=0-(-184/32) and x=0+(-184/32) then you get x=-184/32 and x=184/32. The answer is 184/32 since time can't be negative. To simplify, divide 184 by 32 and you get 5 and 3/4 seconds.